Discuss the methods used to determine safety stock when the stock out cost is known and when the stock out cost is unknown.
Answer #1

Safety stock is common word used by logistic companies to determine the level of extra stock that is maintained to migrate from stockouts i.e a situvation that a product is out of stock.The safety stock calculation is based on

1.Demand of a particular commodity,

2.Lead time

3.Service time(it is defined as the desired probability of meeting the demand)

using this as inputs we have a method called reorder point method to calculate the safety stock.

the safety stock(SS) is calculated using the formula.

SS= z\sigma *\sqrt{E(L) \sigma^2(D) )+(E(D))^2\sigma ^2(L))}

where \alpha is the service level

sigma(l) is defined as standard deviation of mean time

E(l) -mean

then the reorder point is calculated as

ROP=E(L).E(D)+SS